OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

slope 1 y=x-4 x-y=4

OpenStudy (anonymous):

so x-y=4?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

thnx